Lexi has a set of five number cards. The mean of the cards is 5, the mode is 8, the median is 6 and the range is 7. Write down the five numbers on Lexi's cards.
Answer by greenestamps(13367)



The number cards are: A B C D E

Use the given information to determine specific cards.

The median (middle card -- i.e., the 3rd card) is 6; the cards are: A B 6 D E

The mode (most common number) is 8; the cards are: A B 6 8 8

Note that since 8 is the only mode, A and B must be different.

The range (difference between largest and smallest) is 7; the cards are: 1 B 6 8 8

The mean is 5, so the sum is 5*5 = 25.
1 + B + 6 + 8 + 8 = 25
B + 23 = 25
B = 2

ANSWER: The numbers are: 1 2 6 8 8